Gae S.p.A. is an Italian company which fabricates two kinds of cars: model A and model B. The overhead costs are applied based on machine hours. For this accounting period, Vidoni S.p.A. estimated that 41,000 machine hours and incur in a manufacturing overhead equal to $455,000 . The capacity of the company is 50,000 machine hours. The following are the financial data for this accounting period:

Indirect materials $ 35,000

Direct labor 127,500

Indirect manufacturing labour $ 115,000

Factory depreciation $ 285,000

Administrative depreciation $ 75,000

Factory insurance $ 13,750

Administrative insurance $ 8,000

Administrative salaries $ 205,000

General selling expenses $ 23,000

Factory energy costs $ 48,000

Advertising $ 185,000

Fixed revenues $23,000

The information about the two models are the following:

MODEL A MODEL B
Units sold 5000 2700
Price 700 1000
Units produced 5500 2500
Direct materials (per unit) 90 130
Direct labour (per unit) 15 18
Machine hours (per unit) 3 10
Sales commission 5% 8%

Please, keep in consideration that direct labour is FIXED ( according to Italian legislation) and indirect materials are considered variable.

Tasks:

1. Determine the (i) predetermined rate; (ii) effective overhead rate ; (iii) correct and effective fixed cost rate assuming 30% of Factory energy cost is variable.

2. Compute the effective cogs under absorption costing (assume same level of efficiency and production for the past accounting period).

3. Provide COGS income statement.

4. Provide CONTRIBUTION MARGIN income statement.

5. Find the BEP in $ for the company.
Answer #1

Ans: Gae Spa fabricate two car models Model A and Model B

Basic Information

Particular Model A Model B
Sales (unit) 5,000 2,700
Production (unit) 5,500 2,500
Sale Price ($) 700 1,000
Prime Cost (Direct labor + direct material) per unit 105 148
sales commission per unit ($) { (700*5%):(1,000*8%)} 35 80
Machine hour per unit ( hours) 3 10
Machine Hours for production 16,500 25,000

i) a)Predetermined overhead rate = Total estimated manufacturing overhead rate\ Machine hour allocated

= $4,55,000+$75,000+$8,000+$205,000+$23,000+$185,000-$23,000\ 41,000 hours

= $ 22.63

b) Effective overhead rate = Actual overhead cost \ machine hour

Machine hour used = (5,500*3)+(2,500*10)

= 41,500 hrs

Overhead costs = $35,000+$115,000+$285,000+$13,750+$48,000+$75,000+$8,000+$205,000+$23,000+$185,000-$23,000

= $ 969,750

Effective overhead rate = $ 969,750\41,500

= $ 23.36

c) Computation of Fixed Cost

Particulars Amount($)
Indirect manufacturing labour 115,000
Factory depreciation 285,000
Administrative depreciation 75,000
Factory insurance 13,750
Administrative insurance 8,000
Administrative salaries 205,000
General selling expenses 23,000
Factory energy costs ($ 48,000*(1-30%)) 33,600
Advertising    185,000
Total Fixed cost 9,43,350

Fixed cost Rate = Total fixed cost\ Normal machine hours

= $ 9,43,350 \ 41,000

= $ 23

2) Computation of COGS under Absorption costing

Particulars Model A Model B
unit sold 5,000 2,700
Prime cost per Unit 105 148
Prime cost ((5,000*$105):(2,700 *$148)) 525,000 399,600
Fixed cost per unit  
(15,000*$23) 345,000
(27,000*$23) 621,000
Total COGS 870,000 1,020,600

Machine hour for Unit sold

Model A = 5,000* 3 = 15,000 hrs

Model B = 2700* 10 = 27,000 hrs

3) COGS in Income Statement

Particulars Amount ($)
Opening Inventory
Model A  
Model B (200*(148+(23*10))) 75,600
add: Production {( 5,500*(105+(23*3)))+(2,500*(148+(23*10))} 1,902,000
less: closing Stock of model A  ( 500*(105+(23*3))) 87,000
COGS 1,890,600

4) Contribution Margin Income statement

Particulars Model A Model B
Sale Price ($) 700 1,000
less
Prime Cost (Direct labor + direct material) per unit 105 148
Sales commission 35 80
contribution per unit 560 772
contribution 2,800,000 2,084,400

Contribution = $2,800,000 + $2,084,400 - ($48000*30%)- $35000

= $ 4,835,000